OpenStudy (anonymous):

The widget factory is pulling items from each batch to test them for quality control. Twenty-four items are pulled randomly from each batch regardless of the size of the batch. A manager records the percentage and the size of the batch in a log. Which of these could be recorded in the log? 25% of 96 20% of 120 12% of 220 6% of 400 15% of 160 9% of 260

OpenStudy (anonymous):

Convert the percentages and then multiply each by the size of each batch and see if you get 24. 0.25 * 96 = 24 (ANSWER) 0.20 * 120 = 24 (ANSWER) 0.12 * 220 = 26.4 0.06 * 400 = 24 (ANSWER) 0.15 * 160 = 24 (ANSWER) 0.09 * 260 = 23.4 So the answer is the 1st, 2nd, 4th, and 5th ones.

OpenStudy (anonymous):

kk thanks